What is the length of the radius of the circle? (x - 6)2 + (y + 5)2 = 152 please help me
Answer by lwsshak3(11628)

What is the length of the radius of the circle? (x - 6)2 + (y + 5)2 = 152 please help me
..
Standard form of a circle: (x-h)^2+(y-k)^2=r^2, with (h,k) being the (x,y) coordinates of the center and r being tlhe radius.
For given equation,(x-6)^2+(y+5)^2=152
This is a circle with center at (6,-5) and r^2=152
r^2=152
r=√152=12.33..
ans:
radius of given circle=12.33